If ground state of hydrogen atom is chosen as zero potential energy level.The value of total energy in 1st excited state ???

Grade:

2 Answers

vikas askiitian expert
13 years ago

total energy of any level in Hdrogen atom is E = -13.6/n2

E (ground state) = -13.6ev

E(first exited state) = -13.6/4 ev

 

potential of ground state = -27.2ev

if potential of ground state is 0 then we have to add +27.2 , by this method we

can make ground state potential 0 ...

now add this potential to  energy of  each state this will give the final energy of that state ...

E(first exited state) = -13.6/4 + 27.2 = 23.8ev

ankit singh
3 years ago
If potential energy of an electron in a hydrogen atom in first excited state is taken to be zero, kinetic energy (in eV) of an electron in ground state will be. For this to be zero, We must add 6.8 eV.